A vessel and its content have a heat capacity of 500.it is heated using butane burner(the standered enthalpy of combustion of butane is-2877)
a)Calculate the mass of butane needed to give a theoretical temperature rise of 30 degree centigrade
b)Calculate the theoretical temperature rise occurring when 0.50g of butane is burned.
1
Expert's answer
2020-05-19T08:59:47-0400

a)

q = cΔT

500 J / mol K * 30K = 15000 J (15 kJ) - q of calorimeter

1 mol of butane produce 2877 kJ of heat

x = 15/2877 = 0.0052 mol

M (butane) = 58.12

0.0052 * 58.12 = 0.303 g

b)

0.50 / 58.12 = 0.0086 mol

1 mol of butane produce 2877 kJ of heat

x = 0.0086 * 2877 = 24.75 kJ = 24750.5 J

q = cΔT

ΔT = q / c = 24750.5 / 500 = 49.5 K